# Universal spectra of noisy parameterized quantum circuits

### v1

@@ III.2 Synthetic data from CPTP maps
$\mathcal{L}=-i\alpha(I\otimes H-H^{T}\otimes I)+\phi-\frac{1}{2}({\text{Tr}}_{B}(\phi)^{T}\otimes I-I\otimes{\text{Tr}}_{B}(\phi)),$ (S8)
@@
$\begin{split}\rho^{\bm{s}}=\Phi(P_{\bm{s}}\rho_{in}P_{\bm{s}}^{\dagger})\\ p_{j}^{(\bm{s},\bm{b})}=\mathrm{Tr}[E_{j}P_{\bm{b}}\rho^{\bm{s}}P_{\bm{b}}^{\dagger}].\end{split}$ (S9)
@@
In Figure S3, we compare the retrieved spectra (orange dots) with the spectra of the original maps (blue dots). Although the accuracy in matching eigenvalues as a finite set of points on the plane leaves room for further improvement, the procedure is able to capture the spectral support and spectral density quite accurately.
Figure S3: Spectra of original maps (blue dots), Eqs. (S5-S6), and of the retrieved maps (orange dots) for $n=3$ (upper panel) and $n=4$ (lower panel). Parameters of the map are (a) $\alpha=1$, $\beta=0.1$, $r=1$, (b) $\alpha=10^{4}$, $\beta=10^{-3}$, $r=16$, (c) $\alpha=10^{2}$, $\beta=10^{-3}$, $r=16$, (d) $\alpha=1$, $\beta=10^{-2}$, $r=8$.
III.3 Spectral Fitting
Diluted Unitaries (DU) [S3] constitute an ensemble of random maps,
$T_{\text{Diluted}}(\rho)=(1-p)U\rho U^{\dagger}+p\sum_{j=1}^{r}K_{j}\rho K_{j}^{\dagger},$ (S10)
@@ III.3 Spectral Fitting
Figure S4: Histograms of the spectra $\{\lambda_{j}\},\leavevmode\nobreak\ j=2,...,d^{2}$ of the maps $\mathcal{T}_{d^{2}}$ retrieved with three-qubit parameterized circuits of varying depth $\ell$, implemented on ibmq_belem. For each value of $\ell$, we retrieved spectra of ten circuits, obtained by randomly sampling over the rotation angles. Each plot displays a two-dimensional histogram obtained with $63010(2^{6}-1)=$ 630 eigenvalues (eigenvalue $1$ is not included). The red dots represent eigenvalues of a diluted unitary with $p$ and $r$ obtained by minimizing the spectral distance to one (randomly sampled) diluted unitary and one (randomly chosen) realization of the parameterized circuit.
@@
where $p(\lambda;S^{(a)},\sigma)$ is a kernel density estimation over the spectrum $S^{(a)}$, using a ywo-dimensional Gaussian kernel of width $\sigma$. It is given by the sum
$p(\lambda;S^{(a)},\sigma)=\frac{1}{d^{2}}\sum_{i=1}^{d^{2}}\mathcal{N}(\lambda;\lambda^{(a)}_{i},\sigma).$ (S12)
The integral in Eq. S11 can be computed analytically, resulting in the expression
SD($\text{SD}(S^{(a)},S^{(b)})=\frac{1}{d^{4}}\sum_{i=1,j=1}^{d^{2}}\left[\mathcal{N}(\lambda^{(a)}_{i};\lambda^{(a)}_{j},\sigma)+\mathcal{N}(\lambda^{(b)}_{i};\lambda^{(b)}_{j},\sigma)-2\mathcal{N}(\lambda^{(a)}_{i};\lambda^{(b)}_{j},\sigma)\right].$ (S13)
@@
IV Results for parameterized circuit with three qubits
Figure S4, similar to Fig.3 in the main text, presents the spectra retrieved from experimental data collected with three-qubit parameterized circuits of varying depth $\ell$, implemented on ibm_belem. For each depth value, we sampled ten random sets of angles and retrieved the ten corresponding spectra. We calculated two-dimensional histograms using $63010(2^{3\cdot 2}-1)=$ 630 eigenvalues for each value of $\ell$. The histograms are plotted together with the spectra of diluted unitaries with $p$ and $r$ values obtained by minimizing the spectral distance to one (out of ten) of the circuit realizations.

### v2

@@ C.2 Synthetic data from CPTP maps
$\mathcal{L}=-i\alpha(I\otimes H-H^{T}\otimes I)+\phi-\frac{1}{2}({\text{Tr}}_{B}(\phi)^{T}\otimes I-I\otimes{\text{Tr}}_{B}(\phi)),$ (11)
@@
Figure 8: Eigenvalues $\{\lambda_{j}\},\leavevmode\nobreak\ j=2,...,d^{2}$ of the maps $\mathcal{T}_{d^{2}}$ retrieved with three-qubit parameterized circuits of varying depth $\ell$ implemented on ibmq_belem For each depth $l$, ten independent circuits was implemented and retrieved, and the resulting spectra overlaid. The spectra are invariant under the complex conjugation and only the eigenvalues with $\text{Im}\lambda\geqslant 0$ are plotted. The eigenvalues with $\text{Im}\lambda^{\text{DQ}}\leqslant 0$, of the diluted unitaries $\mathcal{DU}_{d}(p,r)$, obtained as spectral approximations of the retrieved maps, are shown together with the corresponding average $p$ and $r$ values. Radii of the outer, $R_{+}$, and inner, $R_{-}$, circles are given by $R_{\pm}=\sqrt{(1-p)^{2}\pm p^{2}/r}$, averaged over the ten realizations. For $\ell=40$, the eigenvalues of the corresponding diluted unitaries filled a disc.
@@
$\begin{split}\rho^{\bm{s}}=\Phi(P_{\bm{s}}\rho_{in}P_{\bm{s}}^{\dagger})\\ p_{j}^{(\bm{s},\bm{b})}=\mathrm{Tr}[E_{j}P_{\bm{b}}\rho^{\bm{s}}P_{\bm{b}}^{\dagger}].\end{split}$ (12)
@@
In Figure 7, we compare the retrieved spectra (orange dots) with the spectra of the original maps (blue dots). Although the accuracy in matching eigenvalues as a finite set of points on the plane leaves room for further improvement, the procedure is able to capture the spectral support and spectral density quite accurately.
C.3 Spectral Fitting
Diluted Unitaries (DU) [20] constitute an ensemble of random maps,
$T_{\text{Diluted}}(\rho)=(1-p)U\rho U^{\dagger}+p\sum_{j=1}^{r}K_{j}\rho K_{j}^{\dagger},$ (13)
@@ C.3 Spectral Fitting
where $p(\lambda;S^{(a)},\sigma)$ is a kernel density estimation over the spectrum $S^{(a)}$, using a two-dimensional Gaussian kernel of width $\sigma$. It is given by the sum
$p(\lambda;S^{(a)},\sigma)=\frac{1}{d^{2}}\sum_{i=1}^{d^{2}}\mathcal{N}(\lambda;\lambda^{(a)}_{i},\sigma).$ (15)
The integral in Eq. 14 can be computed analytically, resulting in the expression
SD($\text{SD}(S^{(a)},S^{(b)})=\frac{1}{d^{4}}\sum_{i=1,j=1}^{d^{2}}\left[\mathcal{N}(\lambda^{(a)}_{i};\lambda^{(a)}_{j},\sigma)+\mathcal{N}(\lambda^{(b)}_{i};\lambda^{(b)}_{j},\sigma)-2\mathcal{N}(\lambda^{(a)}_{i};\lambda^{(b)}_{j},\sigma)\right].$ (16)
@@
Appendix D Results for parameterized circuit with three qubits
Figure 8, similar to Fig. 3 in the main text, presents the spectra retrieved from experimental data collected with three-qubit parameterized circuits of varying depth $\ell$, implemented on ibm_belem. For each depth value, we sampled ten random sets of angles and retrieved the ten corresponding spectra. The spectra are plotted together with the spectra of diluted unitaries with $p$ and $r$ values obtained by minimizing the spectral distance to each of the ten circuits.